# Predictive Coherence and the Moment Hierarchy: Martingale Posteriors for Exchangeable Bernoulli Sequences

### 05 Mar 2026 v2

@@ 1 Introduction
Let $(X_{i})_{i\geq 1}$ be an exchangeable $\{0,1\}$-valued sequence. By de Finetti’s theorem, there exists a unique mixing measure $\Pi$ on $[0,1]$ such that, conditionally on $\theta\sim\Pi$, the $X_{i}$ are i.i.d. $\mathrm{Bern}(\theta)$. The posterior $\Pi(\cdot\mid\mathcal{F}_{n})$ determines all predictive quantities through integration.
Fong, Holmes, and Walker (2023) proposed replacing the prior-times-likelihood mechanism with a single coherence condition: the sequence $(\theta_{n})_{n\geq 0}$ must satisfy
@@
where $\mathcal{F}_{n}=\sigma(X_{1:n})$ and $\mathcal{F}_{0}$ is trivial. By the martingale convergence theorem, $\theta_{n}\to\theta_{\infty}$ a.s., and the law of the directing parameter $\theta_{\infty}$ is a posterior. Condition (1) constrains only the first conditional moment of $\theta_{\infty}$ given $\mathcal{F}_{n}$.
We study the extent to which first-moment coherence determines multi-step predictive quantities. The Bernoulli case is the natural setting: it isolates the structural phenomenon in the simplest exchangeable model.
Theorem (Main results, informal).
The martingale condition identifies the posterior mean; we characterize what additional structure is required for predictive completeness. Specifically: for $k\geq 2$, the mapping $m_{n}\mapsto\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ is set-valued, so first-moment information does not uniquely identify the $k$-step predictive. Under any strictly proper scoring rule, the plug-in rule is strictly dominated by the Bayes predictive whenever the posterior is non-degenerate. A martingale posterior is predictively complete if and only if the conditional law of $\theta_{\infty}$ is uniquely specified.
Goldstein’s (1975a) conditional-prevision programme organises beliefs at the moment level: the analyst commits to selected expectation functionals rather than a full posterior law. Cover’s (1985) pathwise programme accumulates log-evidence at the KL rate, supplying the e-process structure for anytime-valid inference. The martingale posterior of Fong, Holmes, and Walker (2023) imposes a first-moment coherence condition analogous to Goldstein’s prevision commitment; its almost-sure convergence connects it to Cover’s framework. Table 1 summarises these relationships.
Table 1: Four frameworks for posterior belief about an exchangeable sequence, ordered by strength of commitment. Commitment is what the analyst is required to specify; posterior object is what is thereby determined; multi-step uniqueness asks whether block-event predictives are uniquely fixed for every horizon. The non-uniqueness result for the final row is Theorem 6.3.
Framework Commitment Posterior object Multi-step uniqueness
de Finetti coherence Full prior on directing measure Full posterior law Yes (de Finetti’s theorem)
Classical Bayesian updating Likelihood and prior Full posterior law Yes (Bayes’ rule)
Goldstein conditional previsions Chosen expectation functionals (conditional previsions) A finite set of posterior moments Partial: each committed prevision fixes one predictive horizon
Martingale posteriors (this paper) First conditional moment only Terminal-value law No: underdetermined for every horizon beyond one
The first two frameworks inherit full predictive uniqueness from their commitment to a complete posterior law. The Goldstein row makes explicit that committing to finitely many expectation functionals determines only the corresponding predictive horizons. The final row identifies the case studied here: Theorems 6.3 and 10.3 characterise the resulting gap and its closure condition.
De Finetti’s theorem and Sanov’s theorem are dual descriptions of the same sample path. De Finetti provides a directing measure $\mu$ as the almost-sure limit of the empirical measure $L_{n}$; Sanov quantifies the cost of deviating from that limit:
$\mathbb{P}(L_{n}\in A)\;\asymp\;\exp\!\Bigl(-n\inf_{\mu\in A}D_{\mathrm{KL}}(\mu\|\mu_{0})\Bigr).$
Reversing the KL argument (a standard reweighting identity; see e.g. Cover and Thomas 2006, Ch. 12) gives the inverse Sanov identity: posterior mass near $\mu$ concentrates at rate $\exp(-n\,D_{\mathrm{KL}}(L_{n}\|\mu))$, recovering the Bernstein–von Mises theorem as the moderate-deviations limit. The martingale posterior of Fong, Holmes, and Walker (2023) sits at the intersection of both: $(\Pi_{n})$ is a measure-valued martingale whose almost-sure limit $\Pi_{\infty}=\delta_{\mu}$ recovers the de Finetti directing measure, while its contraction rate is governed by $D_{\mathrm{KL}}(\cdot\|\mu_{0})$. Section 2 develops this connection; Section 4 shows that first-moment information does not determine the higher-order structure of $\Pi_{\infty}$ (Theorem 6.3).
Goldstein (1975a) and Goldstein and Wooff (1998) develop the Bayes linear programme in which conditional previsions are the primary inferential object. In a companion paper, Goldstein (1975b) showed that when the posterior mean is a linear function of the sample, a recursive relation uniquely determines all moments of the prior distribution—provided the likelihood moments are known. The present paper establishes that without this likelihood structure, first-moment constraints alone do not identify block-event predictives (Theorem 6.3). Diaconis and Ylvisaker (1979) characterized conjugate priors by linearity of the posterior mean; Condition (1) is the sequential analogue. Polson and Zantedeschi (2025) construct predictive distributions from finite moment constraints; Datta and Polson (2025) show polynomial log-marginals of degree $\geq 3$ cannot correspond to any valid prior. Section 4 identifies which multi-step predictives remain underdetermined under first-moment constraints (Theorem 6.3).
The main results are the moment hierarchy and insufficiency theorem (Theorems 4.1 and 6.3), the plug-in dominance result (Proposition 7.3), and a closure theorem characterizing predictive completeness (Theorem 10.3). Hill’s $A_{(n)}$ rule under the Jeffreys $\mathrm{Beta}(\tfrac{1}{2},\tfrac{1}{2})$ prior serves as a positive example (Section 8). A companion paper (Polson and Zantedeschi, 2026) develops a broader taxonomy of admissibility geometries (Bayes risk, e-process validity, conformal coverage, and Cesàro approachability); the present paper isolates the moment-hierarchy obstruction behind mean-only coherence.
The remainder of the paper is organised as follows. Section 2 develops the sample-path framework and Section 3 sets up the Bernoulli model. Sections 4 through 7 establish the moment hierarchy, insufficiency, and plug-in dominance. Section 8 provides a positive example via Hill’s $A_{(n)}$ rule. Sections 9 and 10 give asymptotic reconciliation and the closure theorem. Section 11 treats optimal stopping, and Section 12 concludes.
@@ 2 Sanov, de Finetti, and the sample-path picture
The probability that $L_{n}$ lies in a measurable set $A$ decays as
$\mathbb{P}(L_{n}\in A)\asymp\exp\!\bigl(-n\inf_{\mu\in A}D_{\mathrm{KL}}(\mu\|\mu_{0})\bigr)$
(Sanov, 1957). The KL divergence is the canonical cost of deviation from the truth.
For an exchangeable sequence $(X_{i})_{i\geq 1}$, de Finetti’s theorem (Theorem 3.2 below) provides a unique prior $\Pi$ on the space of data-generating laws such that, conditionally on $\mu\sim\Pi$, the observations are i.i.d. with law $\mu$. By the strong law, $L_{n}\to\mu$ a.s., so the law of the limit is $\mathcal{L}(L_{\infty})=\Pi$. Each infinite sample path reveals exactly one draw from $\Pi$; the prior itself is not observable from a single path.
@@
where $\mathcal{I}(\theta_{0})$ is the Fisher information matrix. Eichelsbacher and Ganesh (2002) show the posterior satisfies a moderate-deviation principle (MDP) with this quadratic rate, sharpening the Bernstein–von Mises theorem to intermediate scales. For an extension of Sanov’s theorem to Gibbs conditioning and tilted empirical measures, see Léonard and Najim (2002).
An e-variable is a nonnegative random variable $E$ with $\mathbb{E}_{H_{0}}[E]\leq 1$. An e-process $(E_{n})$ is a test martingale under $H_{0}$; Ville’s inequality gives type-I error control at any stopping time (Ville, 1939; Grünwald et al., 2024). The log of the growth-rate-optimal e-process satisfies
$\frac{1}{n}\log E_{n}\;\longrightarrow\;D_{\mathrm{KL}}(\mu\|\mu_{0})\quad\text{a.s.}$ (4)
When $\mu=\mu_{0}$, $E_{n}$ is a martingale and does not drift; when $\mu\neq\mu_{0}$, $E_{n}\to\infty$ at exponential rate $D_{\mathrm{KL}}(\mu\|\mu_{0})>0$. The alternative is discovered from the sample path as $L_{n}$ converges, not specified in advance.
The Bayes factor $E_{n}=\int p(X_{1:n}\mid\mu)\Pi(d\mu)\big/p(X_{1:n}\mid\mu_{0})$ is automatically a test martingale under $H_{0}$. The prior $\Pi$ affects power but not validity.
Table 2 collects the four frameworks. Each describes the information in a single sample path about the directing measure $\mu$.
@@
The martingale posterior of Fong, Holmes, and Walker (2023) sits in the bottom row: it constrains the first moment of $\Pi_{n}$ but not the law. The insufficiency results of Section 6 quantify this gap; the closure theorem (Theorem 10.3) identifies the exact boundary.
@@
Theorem 2.1 (Sanov–Moment Bridge).
@@
    First-moment coherence as linearisation. The martingale condition $\mathbb{E}[\theta\mid\mathcal{F}_{n}]=L_{n}$ constrains only the mean of the posterior (5): it pins the centre of the KL bowl but specifies nothing about the posterior’s spread or shape. First-moment martingale coherence is the linearisation of Bayesian updating: it retains the location of the posterior minimum but discards the curvature encoded in $\nabla^{2}D_{\mathrm{KL}}$ and all higher derivatives.
@@
    Multi-step predictives require curvature. By Corollary 4.2, $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ depends on posterior moments up to order $k$. For $k=2$, the discrepancy between Bayes and plug-in equals the posterior variance $\sigma_{n}^{2}$ (equation (12)), which is governed by the quadratic term in (6). For $k\geq 3$, additional cumulants enter, corresponding to higher derivatives of the Sanov rate.
@@
    Derivative correspondence. Higher-order predictives are sensitive to higher-order features of the KL profile; in the Bernoulli case this is made explicit via the polynomial expansion $(1-\theta)^{k}=\sum_{j}\binom{k}{j}(-\theta)^{j}$ (Theorem 4.1). The moment hierarchy follows from this expansion, not from special structure of the Bernoulli model.
@@
(iii)–(v) follow from (i)–(ii) together with Theorem 4.1 and Corollary 4.2. ∎
@@ 3.1 Exchangeable Bernoulli sequences
Standard; see de Finetti (1937) and Hewitt–Savage (1955). The mixing measure is the law of $\lim_{n\to\infty}S_{n}/n$, identified via the ergodic decomposition of the exchangeable $\sigma$-algebra. ∎
@@ 3.2 Martingale posteriors and the terminal value
Proposition 3.5 (Terminal value).
Under Assumption 3.4, $\theta_{n}\to\theta_{\infty}$ a.s. and in $L^{1}$, and $\mathbb{E}[\theta_{\infty}\mid\mathcal{F}_{n}]=\theta_{n}$ a.s. for all $n$. The random variable $\theta_{\infty}$ is the directing parameter.
@@
Definition 3.7 (Predictive coherence).
@@ 3.3 Martingale versus exchangeability
Proposition 3.8 (Martingale does not imply exchangeability).
@@
Under exchangeability, the posterior mean (7) is a function of $(n,S_{n})$ alone. Let $\theta_{0}=1/2$ and define $\theta_{n}=\theta_{n-1}+2^{-n}(2X_{n}-1)$. By construction $(\theta_{n})$ is a bounded martingale, but $\theta_{n}$ depends on the order of observations, not merely $S_{n}$. Hence it cannot be a posterior mean under exchangeability. ∎
Definition 3.9 (Conditionally identically distributed).
@@
Exchangeability $\Rightarrow$ c.i.d.: the conditional distribution of $X_{n+j}$ given $\mathcal{F}_{n}$ depends only on $(n,S_{n})$ and equals $\mathrm{Bern}(\mathbb{E}[\theta\mid\mathcal{F}_{n}])$ for all $j\geq 1$. For the converse, the martingale in Proposition 3.8 generates a c.i.d. sequence (future observations are conditionally $\mathrm{Bern}(\theta_{n})$) that is not exchangeable; see Berti, Pratelli, and Rigo (2004). ∎
@@
Remark 3.11 (Role of c.i.d. in predictive coherence).
The c.i.d. property guarantees that the one-step conditional distribution $\mathbb{P}(X_{n+1}\in\cdot\mid\mathcal{F}_{n})$ is the same for all future indices $j\geq 1$. This is one-step predictive coherence (Definition 3.7 with $k=1$): knowing $\theta_{n}$ determines $\mathbb{P}(X_{n+j}=1\mid\mathcal{F}_{n})=\theta_{n}$ for every $j$. However, c.i.d. does not imply that future observations are conditionally independent given $\mathcal{F}_{n}$; that requires knowledge of the full conditional law of $\theta_{\infty}$. For $k\geq 2$, the joint predictive $\mathbb{P}(X_{n+1}=x_{1},X_{n+2}=x_{2}\mid\mathcal{F}_{n})$ depends on the posterior variance through (12), which is not determined by the c.i.d. property alone. Thus c.i.d. $\Rightarrow$ 11-step coherence, but c.i.d. $\not\Rightarrow$ $k$-step coherence for $k\geq 2$, paralleling Theorem 6.3 at the process level.

### 01 Apr 2026 v5

# Martingale Posterior Predictive Coherence: Hausdorff Moment Hierarchy

@@ Abstract
Keywords: exchangeability, de Finetti, predictive inference, martingales, admissibility, moment problems.
@@ 1 Introduction
Kreps (1988) posed a simple question: you observe a sequence of coin flips that you believe to be exchangeable; what can you say about the next $k$ outcomes? De Finetti’s (1937) theorem answers in principle: there exists a unique mixing measure $\Pi$ on $[0,1]$ such that every predictive probability is an integral against the posterior $\Pi(\cdot\mid\mathcal{F}_{n})$. But predicting the next flip requires only the posterior mean, while predicting the next run of flips requires the full posterior shape.
For an exchangeable Bernoulli sequence $(X_{i})_{i\geq 1}$, the probability that the next $k$ observations are all zero is the posterior expectation $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$. By the binomial theorem, this depends on all posterior moments up to order $k$, not just the mean. Each additional prediction step demands one additional moment. This is the moment hierarchy.
To see why this matters concretely, consider exchangeable binary failure indicators with posterior mean $\mathbb{E}[\theta\mid\mathcal{F}_{n}]=0.1$. If the posterior is $\delta_{0.1}$, the probability that all four components fail simultaneously is $0.1^{4}=10^{-4}$. If the posterior instead places equal mass on $0.01$ and $0.19$ (same mean), the probability is $\mathbb{E}[\theta^{4}]\approx 6.6\times 10^{-4}$—more than six times larger. The distinction is governed by the fourth posterior moment, which the first moment does not determine.
Fong, Holmes, and Walker (2023) proposed replacing the full prior-times-likelihood machinery with a single coherence condition: the sequence $(\theta_{n})_{n\geq 0}$ is coherent if
@@
where $\mathcal{F}_{n}=\sigma(X_{1:n})$. This is a martingale condition; by the convergence theorem, $\theta_{n}\to\theta_{\infty}$ a.s., and the conditional law of $\theta_{\infty}$ given $\mathcal{F}_{n}$ plays the role of a posterior in the sense that $\mathbb{E}[\theta_{\infty}\mid\mathcal{F}_{n}]=\theta_{n}$. The appeal is parsimony: a single moment constraint, and the rest follows from convergence.
But condition (1) constrains only the first conditional moment of $\theta_{\infty}$ given $\mathcal{F}_{n}$. For one-step prediction this suffices: $\mathbb{P}(X_{n+1}=1\mid\mathcal{F}_{n})=\theta_{n}$. For two-step prediction, $\mathbb{E}[(1-\theta)^{2}\mid\mathcal{F}_{n}]=(1-\theta_{n})^{2}+\mathrm{Var}(\theta\mid\mathcal{F}_{n})$ additionally requires the posterior variance, which the martingale condition does not touch. This paper characterises the full extent of this gap.
The main results of the paper can be summarised informally as follows.
Theorem (Informal summary).
The martingale condition identifies the posterior mean. We characterize what additional structure is required for predictive completeness. Specifically:
  1. (i)
    For $k\geq 2$, the mapping $m_{n}\mapsto\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ is set-valued: first-moment information does not uniquely identify the $k$-step predictive (Theorem 23).
  2. (ii)
    Under any strictly proper scoring rule, the plug-in predictive $(1-\theta_{n})^{k}$ is strictly dominated by the Bayes predictive whenever the posterior is non-degenerate (Proposition 30).
  3. (iii)
    A martingale posterior is predictively complete if and only if the conditional law of $\theta_{\infty}$ is uniquely specified; equivalently, by Hausdorff’s (1921) moment theorem, if and only if all conditional moments are determined (Theorem 47).
Hill’s $A_{(n)}$ rule under the Jeffreys $\mathrm{Beta}(\tfrac{1}{2},\tfrac{1}{2})$ prior serves as a positive example: it specifies the full Beta posterior and thereby achieves predictive completeness at every horizon (Section 8).
The question “how much of the posterior is needed for prediction?” has been approached from several directions. Goldstein’s (1975a) conditional-prevision programme organises beliefs at the moment level: the analyst commits to selected expectation functionals rather than a full posterior law. Goldstein (1975b) showed that when the posterior mean is linear in the data, a recursive relation uniquely determines all prior moments given the likelihood structure; Goldstein and Wooff (1998) developed these ideas into the Bayes linear programme. Diaconis and Ylvisaker (1979) characterised conjugate priors by linearity of the posterior mean; condition (1) is the sequential analogue. The present paper establishes that without this conjugate or likelihood structure, first-moment constraints alone do not identify block-event predictives (Theorem 23), and that under any strictly proper scoring rule, the plug-in predictive is dominated by the Bayes predictive whenever the posterior is non-degenerate (Proposition 30).
[…]
@@ 2 Sanov, de Finetti, and the sample-path picture
The probability that $L_{n}$ lies in a measurable set $A$ decays as $\mathbb{P}(L_{n}\in A)\asymp\exp\bigl(-n\inf_{\mu\in A}D_{\mathrm{KL}}(\mu\|\mu_{0})\bigr)$ (Sanov, 1957). The KL divergence is the canonical cost of deviation from the truth.
For an exchangeable sequence $(X_{i})_{i\geq 1}$, de Finetti’s theorem (Theorem 3 below) provides a unique prior $\Pi$ on the space of data-generating laws such that, conditionally on $\mu\sim\Pi$, the observations are i.i.d. with law $\mu$. By the strong law, $L_{n}\to\mu$ a.s., so the law of the limit is $\mathcal{L}(L_{\infty})=\Pi$. Each infinite sample path reveals exactly one draw from $\Pi$; the prior itself is not observable from a single path.
@@
where $\mathcal{I}(\theta_{0})$ is the Fisher information matrix. Eichelsbacher and Ganesh (2002) show the posterior satisfies a moderate-deviation principle with this quadratic rate, sharpening the Bernstein–von Mises theorem to intermediate scales.
Table 1 collects four sample-path descriptions of the information about the directing measure $\mu$.
@@
The martingale posterior of Fong, Holmes, and Walker (2023) sits in the bottom row: it constrains the first moment of $\Pi_{n}$ but not the law. The insufficiency results of Section 6 quantify this gap; the closure theorem (Theorem 47) identifies the exact boundary.
@@
Theorem 1 (Sanov–Moment Bridge).
@@
    Location versus curvature. The KL representation shows that the posterior (4) is centred near the empirical mean $L_{n}$ in likelihood geometry. $L_{n}$ identifies the likelihood/KL centre, while the posterior mean $m_{n}=\mathbb{E}[\theta\mid\mathcal{F}_{n}]$ is the posterior first moment; the present paper concerns the latter. Knowledge of $m_{n}$ pins the approximate location of the posterior but specifies nothing about its spread or shape: first-moment information discards the curvature encoded in $\nabla^{2}D_{\mathrm{KL}}$ and all higher derivatives.
@@
    Multi-step predictives require curvature. By Corollary 14, $\mathbb{E}[(1-\theta)^{k}\mid\mathcal{F}_{n}]$ depends on posterior moments up to order $k$. For $k=2$, the discrepancy between Bayes and plug-in equals the posterior variance $\sigma_{n}^{2}$ (equation (11)), which is governed by the quadratic term in (5). For $k\geq 3$, additional cumulants enter, corresponding to higher derivatives of the Sanov rate.
@@
    Derivative correspondence. Higher-order predictives are sensitive to higher-order features of the KL profile; in the Bernoulli case this is made explicit via the polynomial expansion $(1-\theta)^{k}=\sum_{j}\binom{k}{j}(-\theta)^{j}$ (Theorem 13). The moment hierarchy follows from this expansion, not from special structure of the Bernoulli model.
@@
(iii)–(v) follow from (i)–(ii) together with Theorem 13 and Corollary 14. ∎
@@ 3.1 Exchangeable Bernoulli sequences
The exchangeable $\sigma$-algebra $\mathcal{E}=\bigcap_{n}\sigma(X_{n+1},X_{n+2},\ldots)$ is a.s. trivial under any ergodic component. By the strong law, $S_{n}/n\to\theta_{\infty}$ a.s., and $\theta_{\infty}$ is $\mathcal{E}$-measurable. The ergodic decomposition expresses $\mathbb{P}$ as a mixture $\mathbb{P}(\cdot)=\int\mathbb{P}_{\theta}(\cdot)\,\mathrm{d}\Pi(\theta)$ where each $\mathbb{P}_{\theta}$ generates i.i.d. $\mathrm{Bern}(\theta)$. Uniqueness of $\Pi$ follows from the fact that the moment sequence $\int_{0}^{1}\theta^{k}\,\mathrm{d}\Pi(\theta)=\lim_{n\to\infty}\mathbb{E}[S_{n}^{(k)}/n^{(k)}]$ is determined by the joint law of the $X_{i}$. See de Finetti (1937) and Hewitt–Savage (1955). ∎
@@ 3.2 Martingale posteriors and the terminal value
Proposition 6 (Terminal value).
Under Assumption 5, $\theta_{n}\to\theta_{\infty}$ a.s. and in $L^{1}$, and $\mathbb{E}[\theta_{\infty}\mid\mathcal{F}_{n}]=\theta_{n}$ a.s. for all $n$. The random variable $\theta_{\infty}$ is the directing parameter.
@@
Definition 8 (Predictive coherence).
@@ 3.3 Martingale versus exchangeability
Proposition 9 (Martingale does not imply exchangeability).
@@
Under exchangeability, the posterior mean (6) is a function of $(n,S_{n})$ alone. We exhibit a $[0,1]$-valued martingale that depends on the order of observations.
Let $(X_{i})_{i\geq 1}$ be i.i.d. $\mathrm{Bern}(1/2)$ and set $c_{i}=2^{-(i+2)}$, so $\sum_{i\geq 1}c_{i}=1/4$. Define
$\theta_{n}=\tfrac{1}{2}+\sum_{i=1}^{n}c_{i}(2X_{i}-1).$
Since $\mathbb{E}[2X_{i}-1\mid\mathcal{F}_{i-1}]=0$, the process $(\theta_{n})$ is an $\mathcal{F}_{n}$-martingale. Because $|\sum_{i=1}^{n}c_{i}(2X_{i}-1)|\leq\sum_{i\geq 1}c_{i}=1/4$, we have $\theta_{n}\in[1/4,\,3/4]\subset[0,1]$ for all $n$. However, $\theta_{n}$ depends on the signed increments $2X_{i}-1$ individually, not merely on $S_{n}$. For example, $\theta_{2}$ takes the value $1/2+c_{1}-c_{2}$ when $(X_{1},X_{2})=(1,0)$ but $1/2-c_{1}+c_{2}$ when $(X_{1},X_{2})=(0,1)$; both have $S_{2}=1$ yet produce different values of $\theta_{2}$. Hence $(\theta_{n})$ cannot be a posterior mean under any exchangeable Bernoulli law. ∎
Definition 10 (Conditionally identically distributed).
@@
Exchangeability $\Rightarrow$ c.i.d.: under exchangeability, the conditional distribution of $X_{n+j}$ given $\mathcal{F}_{n}$ depends only on $(n,S_{n})$ and equals $\mathrm{Bern}(\mathbb{E}[\theta\mid\mathcal{F}_{n}])$ for all $j\geq 1$.
For the converse, let $(\theta_{n})$ be the bounded martingale from Proposition 9, and define a predictive scheme by $X_{n+1}\mid\mathcal{F}_{n}\sim\mathrm{Bern}(\theta_{n})$. We claim this sequence is c.i.d. For each $j\geq 1$, the conditional distribution of $X_{n+j}$ given $\mathcal{F}_{n}$ is Bernoulli, so its law is determined by its mean. By the martingale property, $\mathbb{E}[X_{n+j}\mid\mathcal{F}_{n}]=\mathbb{E}[\theta_{n+j-1}\mid\mathcal{F}_{n}]=\theta_{n}$, so $X_{n+j}\mid\mathcal{F}_{n}\sim\mathrm{Bern}(\theta_{n})$ for all $j\geq 1$. Since the conditional laws are identical (not merely the conditional means), the sequence is c.i.d. (Definition 10). But the sequence is not exchangeable, because $\theta_{n}$ depends on the order of observations rather than on $S_{n}$ alone (Proposition 9). See also Berti, Pratelli, and Rigo (2004). ∎
@@
Remark 12 (Role of c.i.d. in predictive coherence).
The c.i.d. property guarantees that the one-step conditional distribution $\mathbb{P}(X_{n+1}\in\cdot\mid\mathcal{F}_{n})$ is the same for all future indices $j\geq 1$. This is one-step predictive coherence (Definition 8 with $k=1$): knowing $\theta_{n}$ determines $\mathbb{P}(X_{n+j}=1\mid\mathcal{F}_{n})=\theta_{n}$ for every $j$. However, c.i.d. does not imply that future observations are conditionally independent given $\mathcal{F}_{n}$; that requires knowledge of the full conditional law of $\theta_{\infty}$. For $k\geq 2$, the joint predictive $\mathbb{P}(X_{n+1}=x_{1},X_{n+2}=x_{2}\mid\mathcal{F}_{n})$ depends on the posterior variance through (11), which is not determined by the c.i.d. property alone. Thus c.i.d. $\Rightarrow$ 11-step coherence, but c.i.d. $\not\Rightarrow$ $k$-step coherence for $k\geq 2$, paralleling Theorem 23 at the process level.